[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue Nov 07, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Tue Nov 7 11:30:37 UTC 2017


HSU Shift Log for Tue Nov 07 2017 06:30:31 GMT-0500 (EST)

Shift: H

Forecaster: E Blake

Technical Issues Summary:

ATCF issues:

N-AWIPS issues: Bonefish is hung up or loading extremely slowly (25 min)  
this morning. Not sure if anything was done yesterday but it is practically  
unusable t

AWIPS-2/NCP issues:

Other Techincal issues:

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20171107/6885a868/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list